# Impact of memory-burdened black holes on primordial gravitational waves in light of Pulsar Timing Array

### v1

@@ V Conclusions
Notably, we show that, due to the memory burden effect, PBHs as light as $10^{2}–10^{3}$ g can leave detectable imprints on BGWs at higher frequencies while still being compatible with PTA data (see Figs. 3 and 5). Additionally, we observe that the PTA fit is relatively insensitive to the parameter $q$ (see, Figs. 2 and 4), representing the fractional mass retained in PBHs before entering the quantum evaporation phase. However, for smaller $q$ values ($q<0.5$), memory-burdened PBHs may produce a unique kink-like feature in the GW spectrum (see Fig. 5).
@@ Appendix A The memory-burdened PBH evolution sensitivity to $q$
due to the second EMD caused by the memory burden effect, where $T_{\rm dom2}\simeq\left(\frac{q}{1-q}\right)T_{\rm mb}$ taking $q<0.5$. In this case, the transfer functions are given by
@@ Appendix B Fitting PTA data with semiclassical PBHs
In Fig. 7, we show the posteriors of inflationary and PBH parameters for semiclassical PBHs (without the memory burden effect). It is clear that, although BGWs diluted by standard PBH domination fit the PTA data, only “heavier” PBHs are allowed, which is in sharp contrast to memory-burdened case.

### v2

@@ V Conclusions
Notably, we show that, due to the memory burden effect, PBHs as light as $10^{2}–10^{3}$ g can leave detectable imprints on BGWs at higher frequencies while still being compatible with PTA data (see Figs. 3 and 5). Additionally, we observe that the PTA fit is relatively insensitive to the parameter $q$ (see Figs. 2 and 4), representing the fractional mass retained in PBHs before entering the quantum evaporation phase. However, for smaller $q$ values ($q<0.5$), memory-burdened PBHs may produce a unique kink-like feature in the GW spectrum (see Fig. 5).
@@ Appendix A The memory-burdened PBH evolution sensitivity to $q$
due to the second EMD caused by the memory burden effect, where $T_{\rm dom2}\simeq\left(\frac{q}{1-q}\right)T_{\rm mb}$, taking $q<0.5$. In this case, the transfer functions are given by
@@ Appendix B Fitting PTA data with semiclassical PBHs
In Fig. 7, we show the posteriors of inflationary and PBH parameters for semiclassical PBHs (without the memory burden effect). It is clear that, although BGWs diluted by standard PBH domination fit the PTA data, only “heavier” PBHs are allowed, which is in sharp contrast to the memory-burdened case.
Appendix C Impact of finite running of the tensor spectral index
To maintain a model-independent approach in our analysis, we examine the impact of a finite running of the tensor spectral index, $\alpha_{T}=dn_{T}/d\ln k$, on the SGWB spectrum. In this scenario, the inflationary tensor power spectrum, $\mathcal{P}_{T}(k)$, is modified by an additional multiplicative factor given by
$\mathcal{P}_{T}(k)\propto\left(\frac{k}{k_{*}}\right)^{\alpha_{T}\ln(f/f_{*})}$ (43)
where $f_{*}$ denotes the frequency at the pivot scale. Notably, for higher frequencies, this correction can significantly influence the SGWB spectrum. Depending on the sign and magnitude of $\alpha_{T}$, the running can either enhance or suppress the gravitational wave signal at observationally relevant frequencies, introducing an additional source of uncertainty during data interpretation. For a detailed discussion on the effects of running in realistic inflationary models, we refer the readers to ref. [110].
Figure 8: Impact of the finite running of the tensor spectral index, $\alpha_{T}=$ $2\times 10^{-3}$(dashed), 00 (solid), and $-2\times 10^{-3}$ (dotted), on the BGW spectrum, in case of the NANOGrav-15yrs (left panel) and IPTA-DR2 (right panel) best-fits, with a fixed $q=0.5$ (see Tab. 1).
In Fig. 8, we present the BGW spectra for different values of $\alpha_{T}$. It is evident that a nonzero $\alpha_{T}$ can have a substantial impact on the SGWB spectrum over a wide range of frequencies, introducing an additional parameter that significantly influences the PBH parameter space. However, a detailed investigation of its effects on the PBH parameter space is beyond the scope of this work, as it would require a model-dependent analysis.